# Reflected generalized BDSDEs driven by non-homogeneous Lévy processes and obstacle problems for stochastic integro-PDEs with nonlinear Neumann boundary conditions

### v2

@@ 3.2 A priori estimates and uniqueness
$\begin{split}&\mathbb{E}\Phi^{\theta,\mu}_{t}\left|\bar{Y}_{t}\right|^{2}+\theta\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|\bar{Y}_{s}\right|^{2}dV_{s}+\mu\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|\bar{Y}_{s}\right|^{2}d\kappa_{s}+\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left\|\bar{Z}_{s}\gamma_{s}\right\|^{2}_{\ell^{2}}ds\\ &\leq 2\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\bar{Y}_{s}\left(f(s,Y_{s},Z_{s})-f(s,Y^{\prime}_{s},Z^{\prime}_{s})\right)ds+2\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\bar{Y}_{s}\left(h(s,Y_{s})-h(s,Y^{\prime}_{s})\right)d\kappa_{s}\\ &\quad+\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|g(s,Y_{s},Z_{s})-g(s,Y^{\prime}_{s},Z^{\prime}_{s})\right|^{2}ds\\ &\leq 2\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|\bar{Y}_{s}\right|^{2}dV_{s}+\left(\alpha+\frac{1}{2}\right)\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left\|\bar{Z}_{s}\gamma_{s}\right\|^{2}_{\ell^{2}}ds.\end{split}$
@@
Therefore, we have $Z=Z^{\prime}$. On the other hand, by applying the Burkholder-Davis-Gundy inequality and performing similar calculations as those in Proposition 8, together with (3.15), we derive that $\left\|\bar{Y}\right\|^{2}_{\mathcal{S}^{2}_{\theta,\mu}}=0$ for any $(\theta,\mu)\in(2,+\infty)\times(0,+\infty)$. Thus, $Y=Y^{\prime}$ (in the indistinguishable sense). Finally, from (1.5)-(i), we have
@@ 3.3 Existence
3.3.1 Case where generators $f$ do not depend on $z$, and $g$ does not depend on $(y,z)$
@@
From (H-M)-(ii’)-(iii’) and [37, Annex B, p. 524], it follows that for every, $\left(\omega,t,y\right)\in\Omega\times[0,T]\times\mathbb{R}$ and $\delta>0$, there exists a unique $J_{\delta}^{\mathfrak{f}}=J_{\delta}^{\mathfrak{f}}(\omega,t,y)$, $J_{\delta}^{h}=J_{\delta}^{h}(\omega,t,y)\in\mathbb{R}$ such that
@@
From [37, Annex B, Proposition 6.7], recall that for every $y\in\mathbb{R}$, the processes $\mathfrak{f}_{\delta}(\cdot,\cdot,y)$ and $h_{\delta}(\cdot,\cdot,y):\Omega\times[0,T]\rightarrow\mathbb{R}$ are jointly measurable and $\mathcal{F}_{t}$-measurable for each $t\in[0,T]$. Furthermore, it holds that:
@@
The remainder of the proof relies on the uniform estimation (3.26) and the Burkholder-Davis-Gundy inequality. Using a similar computation as that in Proposition 8, we can easily derive that, for any $\theta,\mu>0$,
@@
Now, using the Burkholder-Davis-Gundy inequality and following a similar computation as in Proposition 8 (or 9), along with the convergence result (3.28), we can derive that
@@
$\mathfrak{f}_{\delta}\left(s,Y^{\delta}_{s}\right)=\mathfrak{f}\left(s,Y^{\delta}_{s}+\delta\mathfrak{f}_{\delta}\left(s,Y^{\delta}_{s}\right)\right)~\mbox{ and }~h_{\delta}\left(s,Y^{\delta}_{s}\right)=h\left(s,Y^{\delta}_{s}+\delta h_{\delta}\left(s,Y^{\delta}_{s}\right)\right).$
@@
Using the Burkholder-Davis-Gundy inequality and (3.28), we derive that $\lim_{\delta\rightarrow 0^{+}}\sum_{k=1}^{d}\int_{\cdot}^{T}Z^{\delta,(k)}_{s}\,dH^{(k)}_{s}=\sum_{k=1}^{d}\int_{\cdot}^{T}Z^{(k)}_{s}\,dH^{(k)}_{s}$ in $\mathcal{S}^{2}_{\theta,\mu}$. Then, taking the limit term by term in $\mathbb{L}^{2}$ in (3.19) along with the obtained convergence result, we get
@@
Part II: Beyond the boundedness condition (3.17)
@@
By Part I, there exists a unique process $\left(Y^{\mathsf{p}},Z^{\mathsf{p}},K^{\mathsf{p}}\right)\in\mathcal{B}^{2}_{\theta,\mu}$ such that
@@
Now let $\mathsf{q}>\mathsf{p}$. By using Itô’s formula, the Skorokhod condition in (3.33), the Burkholder-Davis-Gundy inequality, and familiar computations, we end up with
@@
$K^{\mathsf{p}}_{t}=Y^{\mathsf{p}}_{0}-Y^{\mathsf{p}}_{t}-\int_{0}^{t}{\mathfrak{f}}_{\mathsf{p}}(s,Y^{\mathsf{p}}_{s})ds-\int_{0}^{t}{h}_{\mathsf{p}}(s,Y^{\mathsf{p}}_{s})d\kappa_{s}-\int_{0}^{t}\mathfrak{g}(s)\overleftarrow{dB}_{s}+\sum_{k=1}^{d}\int_{0}^{t}Z^{\mathsf{p},(k)}_{s}dH^{(k)}_{s}$.
@@ 3.3.2 General case
Here, the drivers are taken in full generality, with $f$ depending on $z$ and $g$ on $(y,z)$. We now present the section’s main result.
@@
Choosing $\varepsilon_{1}>0$ such that $\varepsilon_{1}<1-\alpha$, we deduce that $\{Y^{n},Z^{n}\}_{n\geq 0}$ is a Cauchy sequence in the Banach space $\mathcal{H}^{2,\mathcal{Q}}_{\theta,\mu}\times\mathcal{H}^{2,\ell^{2}}_{\theta,\mu}$ for any $\theta\geq\theta_{0}$ and $\mu>0$. It remains to show that the sequence $\{Y^{n}\}_{n\geq 0}$ is a Cauchy sequence in the Banach space $\mathcal{S}^{2}_{\theta,\mu}$. To this end, for any two integers $n,m\geq 0$, we set $\mathcal{R}^{n,m}=\mathcal{R}^{n}-\mathcal{R}^{m}$ for $\mathcal{R}^{n}\in\{Y^{n},Z^{n},K^{n}\}$. Then, we simply apply Corollary 7 to $\Phi^{\theta,\mu}_{\tau}|Y^{n+1,m+1}_{\tau}|^{2}$ along with the Burkholder-Davis-Gundy inequality, and we follow a similar argument as in Theorem 10, which yields
@@
$\begin{split}\mathbb{E}\left(\int_{0}^{T}f(s,Y^{n+1}_{s},Z^{n}_{s})ds\right)^{2}+\mathbb{E}\left(\int_{0}^{T}h(s,Y^{n}_{s})d\kappa_{s}\right)^{2}\leq\mathfrak{c}_{\theta,\mu,\zeta}.\end{split}$
@@
By the Burkholder-Davis-Gundy inequality, (H-M)-(vii), and (3.43), we have
@@
Using once more (3.43) and the Burkholder-Davis-Gundy inequality, which yields to
@@
Finally, using the same logic as in Part I of the proof of Theorem 10, we can easily see that the process $(Y,Z,K)$ solves the RGBDSDE-NL (1.5). ∎
@@ 4 Application to an obstacle problem of stochastic IPDE with nonlinear Neumann boundary conditions
This section establishes a probabilistic formula for a stochastic viscosity solution to an obstacle problem involving stochastic integral-partial differential equations with nonlinear Neumann boundary conditions (SIPDE-NBC) of parabolic type, utilizing the RGBDSDE-NL (1.5) under specific assumptions. Our technique entails initially determining the outcome for a SIPDE-NBC. The primary conclusion subsequently arises by modifying the reasons articulated in [1].
For the purpose of conducting a rigorous analysis, we assume that the inhomogeneous Lévy process $L$ lacks a Brownian component and features bounded jumps, i.e.
$L_{t}=\int_{0}^{t}{b}_{s}\,ds+\int_{0}^{t}\int_{|e|<1}e\,\tilde{\mu}(ds,de),\quad t\in[0,T].$ (4.47)
Define $\mathbf{F}^{B}\triangleq\left\{\mathcal{F}_{t,T}^{B}\right\}_{0\leq t\leq T}$. We designate by $\mathcal{M}_{0,T}^{B}$ the set of all $\mathbf{F}^{B}$-stopping times $\tau$ satisfying $0\leq\tau\leq T$, $\mathbb{P}$-almost surely. We delineate the subsequent vector spaces of functions for arbitrary Euclidean spaces $D$ and $D_{1}$:
  • $\mathcal{C}^{k_{1},k_{2}}\left([0,T]\times D;D_{1}\right)$ is the space of all $D_{1}$-valued functions defined on $0,T]\times D$ which are $k_{1}$ times continuously differentiable in $t$ and $k_{2}$ times continuously differentiable in $x$, and $\mathcal{C}^{k_{1},k_{2}}_{b}\left([0,T]\times D;D_{1}\right)$ denotes the subspace of $\mathcal{C}^{k_{1},k_{2}}\left([0,T]\times D;D_{1}\right)$ in which all functions have uniformly bounded partial derivatives;
  • $\mathcal{C}^{k_{1},k_{2}}(\mathcal{G},[0,T]\times D;D_{1})$ (resp. $\mathcal{C}_{b}^{k_{1},k_{2}}\left(\mathcal{G},[0,T]\times D;D_{1}\right)$), for any sub-$\sigma$-field $\mathcal{G}\subseteq\mathcal{F}_{T}^{B}$, represents the space of all $\mathcal{C}^{k_{1},k_{2}}\left([0,T]\times D;D_{1}\right)$ (resp. $\mathcal{C}_{b}^{k_{1},k_{2}}\left([0,T]\times D;D_{1}\right)$) random variables valued in $\mathbb{R}$ that are measurable with respect to $\mathcal{G}\otimes\mathcal{B}([0,T]\times D)$.
  • $\mathcal{C}^{k_{1},k_{2}}\left(\mathbf{F}^{B},[0,T]\times D;D_{1}\right)$ (resp. $\mathcal{C}_{b}^{k_{1},k_{2}}\left(\mathbf{F}^{B},[0,T]\times D;D_{1}\right)$) is the space of all random fields $\bar{\alpha}\in\mathcal{C}^{k_{1},k_{2}}\left(\mathcal{F}_{T}^{B},[0,T]\times D;D_{1}\right)$ (resp. $\in\mathcal{C}_{b}^{k_{1},k_{2}}\left(\mathcal{F}_{T}^{B},[0,T]\times D;D_{1}\right)$), such that for fixed $x\in D$, the mapping $(t,w)\rightarrow\bar{\alpha}(t,\omega,x)$ is $\mathbf{F}^{B}$-progressively measurable.
  • $L^{p}(\mathcal{G};D)$, for any sub-$\sigma$-field $\mathcal{G}\subseteq\mathcal{F}_{T}^{B}$ and real number $p\geq 0,$ stands for all $D$-valued $\mathcal{G}$-measurable random variables $\xi$ such that $\mathbb{E}|\xi|^{p}<\infty$.

### v3

@@ 3.2 A priori estimates and uniqueness
$\begin{split}&\mathbb{E}\Phi^{\theta,\mu}_{t}\left|\bar{Y}_{t}\right|^{2}+\theta\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|\bar{Y}_{s}\right|^{2}dV_{s}+\mu\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|\bar{Y}_{s}\right|^{2}d\kappa_{s}+\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left\|\bar{Z}_{s}\gamma_{s}\right\|^{2}_{\ell^{2}}ds\\ &\leq 2\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\bar{Y}_{s}\left(f(s,Y_{s},Z_{s})-f(s,Y^{\prime}_{s},Z^{\prime}_{s})\right)ds+2\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\bar{Y}_{s}\left(h(s,Y_{s})-h(s,Y^{\prime}_{s})\right)d\kappa_{s}\\ &\quad+\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|g(s,Y_{s},Z_{s})-g(s,Y^{\prime}_{s},Z^{\prime}_{s})\right|^{2}ds\\ &\leq 2\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|\bar{Y}_{s}\right|^{2}dV_{s}+\left(\alpha+\frac{1}{2}\right)\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left\|\bar{Z}_{s}\gamma_{s}\right\|^{2}_{\ell^{2}}ds\end{split}$
@@
Therefore, we have $Z=Z^{\prime}$. On the other hand, by applying the B-D-G inequality and performing similar calculations as those in Proposition 9, together with (3.15), we derive that $\left\|\bar{Y}\right\|^{2}_{\mathcal{S}^{2}_{\theta,\mu}}=0$ for any $(\theta,\mu)\in(2,+\infty)\times(0,+\infty)$. Thus, $Y=Y^{\prime}$. Finally, from (1.5)-(i), we have
@@ 3.3 Existence
3.3.1 Case where generators $f$ do not depend on $z$ and $g$ does not depend on $(y,z)$
@@
From (H-M)-(ii’)-(iii’) and [36, Annex B, p. 524], it follows that for every, $\left(\omega,t,y\right)\in\Omega\times[0,T]\times\mathbb{R}$ and $\delta>0$, there exists a unique $J_{\delta}^{\mathfrak{f}}=J_{\delta}^{\mathfrak{f}}(\omega,t,y)$, $J_{\delta}^{h}=J_{\delta}^{h}(\omega,t,y)\in\mathbb{R}$ such that
@@
From [36, Annex B, Proposition 6.7], recall that for every $y\in\mathbb{R}$, the processes $\mathfrak{f}_{\delta}(\cdot,\cdot,y)$ and $h_{\delta}(\cdot,\cdot,y):\Omega\times[0,T]\rightarrow\mathbb{R}$ are jointly measurable and $\mathcal{F}_{t}$-measurable for each $t\in[0,T]$. Furthermore, it holds that:
@@
The remainder of the proof relies on the uniform estimation (3.26) and the B-D-G inequality. Using a similar computation as that in Proposition 9, we can easily derive that, for any $\theta,\mu>0$,
@@
Now, using the B-D-G inequality and following a similar computation as in Proposition 9 (or 10), along with the convergence result (3.28), we can derive that
@@
$\mathfrak{f}_{\delta}\left(s,Y^{\delta}_{s}\right)=\mathfrak{f}\left(s,Y^{\delta}_{s}+\delta\mathfrak{f}_{\delta}\left(s,Y^{\delta}_{s}\right)\right)~\mbox{ and }~h_{\delta}\left(s,Y^{\delta}_{s}\right)=h\left(s,Y^{\delta}_{s}+\delta h_{\delta}\left(s,Y^{\delta}_{s}\right)\right)$
@@
Using the B-D-G inequality and (3.28), we derive that $\lim_{\delta\rightarrow 0^{+}}\sum_{k=1}^{d}\int_{\cdot}^{T}Z^{\delta,(k)}_{s}\,dH^{(k)}_{s}=\sum_{k=1}^{d}\int_{\cdot}^{T}Z^{(k)}_{s}\,dH^{(k)}_{s}$ in $\mathcal{S}^{2}_{\theta,\mu}$. Then, taking the limit term by term in $\mathbb{L}^{2}$ in (3.19) along with the obtained convergence result, we get
@@
Step 2: General case
@@
By Part 1, there exists a unique process $\left(Y^{\mathsf{p}},Z^{\mathsf{p}},K^{\mathsf{p}}\right)\in\mathcal{B}^{2}_{\theta,\mu}$ such that
@@
Now let $\mathsf{q}>\mathsf{p}$. By using Itô’s formula, the Skorokhod condition in (3.33), the B-D-G inequality, and familiar computations, we end up with
@@
$K^{\mathsf{p}}_{t}=Y^{\mathsf{p}}_{0}-Y^{\mathsf{p}}_{t}-\int_{0}^{t}{\mathfrak{f}}_{\mathsf{p}}(s,Y^{\mathsf{p}}_{s})ds-\int_{0}^{t}{h}_{\mathsf{p}}(s,Y^{\mathsf{p}}_{s})d\kappa_{s}-\int_{0}^{t}\mathfrak{g}(s)\overleftarrow{dB}_{s}+\sum_{k=1}^{d}\int_{0}^{t}Z^{\mathsf{p},(k)}_{s}dH^{(k)}_{s}$
@@ 3.3.2 General case
We now provide the main finding of this section.
@@
Choosing $\varepsilon_{1}>0$ such that $\varepsilon_{1}<1-\alpha$, we deduce that $\{Y^{n},Z^{n}\}_{n\geq 0}$ is a Cauchy sequence in the Banach space $\mathcal{H}^{2,\mathcal{Q}}_{\theta,\mu}\times\mathcal{H}^{2,\ell^{2}}_{\theta,\mu}$ for any $\theta\geq\theta_{0}$ and $\mu>0$. It remains to show that the sequence $\{Y^{n}\}_{n\geq 0}$ is a Cauchy sequence in the Banach space $\mathcal{S}^{2}_{\theta,\mu}$. To this end, for any two integers $n,m\geq 0$, we set $\mathcal{R}^{n,m}=\mathcal{R}^{n}-\mathcal{R}^{m}$ for $\mathcal{R}^{n}\in\{Y^{n},Z^{n},K^{n}\}$. Then, we simply apply Corollary 8 to $\Phi^{\theta,\mu}_{\tau}|Y^{n+1,m+1}_{\tau}|^{2}$ along with the B-D-G inequality, and we follow a similar argument as in Theorem 11, which yields
@@
$\begin{split}\mathbb{E}\left(\int_{0}^{T}f(s,Y^{n+1}_{s},Z^{n}_{s})ds\right)^{2}+\mathbb{E}\left(\int_{0}^{T}h(s,Y^{n}_{s})d\kappa_{s}\right)^{2}.\leq\mathfrak{c}_{\theta,\mu,\zeta}.\end{split}$
@@
By the B-D-G inequality, (H-M)-(vii), and (3.43), we have
@@
Using once more (3.43) and the B-D-G inequality, which yields to
@@
Finally, using the same logic as in Part 1 of the proof of Theorem 11, we can easily see that the process $(Y,Z,K)$ solves the RGBDSDE-NL (1.5). ∎
@@ 4 Application to an obstacle problem of stochastic IPDE with nonlinear Neumann boundary conditions
In this section, with the assistance of the RGBDSDE-NL (1.5), we give a probabilistic formula for a stochastic viscosity solution for an obstacle problem of stochastic integral-partial differential equations with nonlinear Neumann boundary conditions (SIPDE-NBC) of parabolic type under a special assumptions.
Our strategy involves first establishing the result for SIPDE-NBC. The main result then follows by adapting the arguments presented in [1].
To facilitate a rigorous analysis, we assume without loss of generality that the inhomogeneous Lévy process $L$ is devoid of a Brownian component and possesses bounded jumps, i.e.
$L_{t}=\int_{0}^{t}{b}_{s}\,ds+\int_{0}^{t}\int_{|z|<1}z\,\tilde{\mu}(ds,dz),\ t\leq T.$ (4.47)
Let $\mathbf{F}^{B}\triangleq\left\{\mathcal{F}_{t,T}^{B}\right\}_{0\leq t\leq T}$. By $\mathcal{M}_{0,T}^{B}$ we will denote all the $\mathbf{F}^{B}$-stopping times $\tau$ such that $0\leq\tau\leq T,\mathbb{P}$-a.s.. For generic Euclidean spaces $E$ and $E_{1}$ we introduce the following vector spaces of functions:
  • $\mathcal{C}^{k_{1},k_{2}}\left([0,T]\times E;E_{1}\right)$ stands for the space of all $E_{1}$-valued functions defined on $0,T]\times E$ which are $k_{1}$ times continuously differentiable in $t$ and $k_{2}$ times continuously differentiable in $x$, and $\mathcal{C}^{k_{1},k_{2}}_{b}\left([0,T]\times E;E_{1}\right)$ denotes the subspace of $\mathcal{C}^{k_{1},k_{2}}\left([0,T]\times E;E_{1}\right)$ in which all functions have uniformly bounded partial derivatives.
  • For any sub-$\sigma$-field $\mathcal{G}\subseteq\mathcal{F}_{T}^{B},\mathcal{C}^{k_{1},k_{2}}\left(\mathcal{G},[0,T]\times E;E_{1}\right)\left(\right$.or $\left.\mathcal{C}_{b}^{k_{1},k_{2}}\left(\mathcal{G},[0,T]\times E;E_{1}\right)\right)$ denotes the space of all $\mathcal{C}^{k_{1},k_{2}}\left([0,T]\times E;E_{1}\right)\left(\right$.or $\mathcal{C}_{b}^{k_{1},k_{2}}\left([0,T]\times E;E_{1}\right)$-valued random variables that are $\mathcal{G}\otimes\mathcal{B}([0,T]\times E)$-measurable.
  • $\mathcal{C}^{k_{1},k_{2}}\left(\mathbf{F}^{B},[0,T]\times E;E_{1}\right)$ $\left(\text{or}\ \mathcal{C}_{b}^{k_{1},k_{2}}\left(\mathbf{F}^{B},[0,T]\times E;E_{1}\right)\right)$ is the space of all random fields $\alpha\in\mathcal{C}^{k_{1},k_{2}}\left(\mathcal{F}_{T}^{B},[0,T]\times E;E_{1}\right)$ $\left(\right$. or $\left.\mathcal{C}_{b}^{k_{1},k_{2}}\left(\mathcal{F}_{T}^{B},[0,T]\times E;E_{1}\right)\right)$, such that for fixed $x\in E$, the mapping $(t,w)\rightarrow\alpha(t,\omega,x)$ is $\mathbf{F}^{B}$-progressively measurable.
  • For any sub-$\sigma$-field $\mathcal{G}\subseteq\mathcal{F}_{T}^{B}$ and real number $p\geq 0,L^{p}(\mathcal{G};E)$ stands for all $E$-valued $\mathcal{G}$-measurable random variables $\xi$ such that $\mathbb{E}|\xi|^{p}<\infty$.